OpenStudy (anonymous):

48 is what percent more than 40

OpenStudy (anonymous):

20% more

OpenStudy (anonymous):

8/40=.2 20 percent more. Just a guess. Been awhile since I did this.

OpenStudy (anonymous):

\[ \frac{48-40}{40} \times 100 = 20 \]